OpenStudy (clamin):

PLEASE HELP!! Calculate the number that you need to add to each side of the equation w^2 - 3w= 350 to create a perfect square trinomial.

OpenStudy (anonymous):

two steps half of 3 is ?

OpenStudy (mathmale):

@clamin: Haven't you posted several "complete the square" problems just recently? If so, what did you learn from working them that you could apply to solving the present problem? Please put into words what YOU have learned fromthis process.

OpenStudy (wolf1728):

To complete the square, the "squared" term of the equation must be 1. The aw^2 + bw terms must be on one side of the equation and the "c" term must be on the other. Luckily for you, the equation is already in this form.

OpenStudy (anonymous):

The form for completing the square is this: \[x ^{2}+\frac{ b }{ 2a }x+\left( \frac{ b }{ 2a } \right)^{2}=0\] use this form to find the squared term. This equation comes from the general form of a quadratic which is: \[ax^{2}+bx+c=0\]

OpenStudy (wolf1728):

w^2 - 3w= 350 Basically the "b" term is 3w. Take the coefficient "3", divide it by 2 and then square it. This equals 2.25. Add this to both sides of the equation: w^2 - 3w + 2.25 = 352.25

OpenStudy (clamin):

@wolf1728 how do i factor this?? is it like this?? w^2 -3w +2.25 - 352.25=352.25 - 352.25 w^2-3w-350=0

OpenStudy (mathmale):

If you do that, you've simply returned to your starting point. Please, review the concept "completing the square." You are asked to re-write the given equation w^2 - 3w= 350 to create a perfect square trinomial on the left side. Then, you should re-write that perfect square trinomial as